OpenStudy (karon86):

3x+4y= -23 2y-x= -19

OpenStudy (sand-lock53):

What's the problem?

OpenStudy (karon86):

A. (-5, -2) B. (3, -8) C. (4, 6) D. (9, -6)

OpenStudy (sweetburger):

solve for x and substitute

OpenStudy (karon86):

okay

OpenStudy (karon86):

If i multiply by -2 will 2x - 4y= 38 work for the top equation?

OpenStudy (sweetburger):

yes it would

OpenStudy (karon86):

-3 = x

OpenStudy (sweetburger):

that is indeed one way to go about this question

OpenStudy (sweetburger):

I did not get x=-3

OpenStudy (karon86):

Alright let me resolve it

OpenStudy (karon86):

So I have 2x - 4y= 38 after I multiply 2 to -x+2y= - 19

OpenStudy (sweetburger):

I'm sorry I read the equation wrong I think your x =-3 is in fact correct.

OpenStudy (karon86):

Then why isn't it one of the answer choice? :(

OpenStudy (sweetburger):

x=+3

OpenStudy (karon86):

how?

OpenStudy (sweetburger):

5x=15 so x=3

OpenStudy (karon86):

how'd you get -15

OpenStudy (karon86):

Oh XDDDD

OpenStudy (karon86):

Found out

OpenStudy (sweetburger):

alright

OpenStudy (karon86):

I got 3, -8

OpenStudy (karon86):

Is that correct?

OpenStudy (oryzae):

Yup! Good job.

OpenStudy (karon86):

Yay!

OpenStudy (karon86):

Thank you guys so much